CHA

lu evenuc T-account and the additional carned revenue in the Service Revenue T-account. the ending balances of the T-accounts under both approaches, Are they ance the same? 1,150. Problems Group A rnalizing adjusting entries and subsequent journal entries 3A Jo Learning Objective 3 Laug hter Landscaping has collected the following data for the December 31 adjusting entries: . Each Friday, Laughter pays employees for the current weeks work. The amount of the weekly payroll is $8,000 for a five-day workweek. This year, December 31 falls on a Tuesday. Laughter will pay its employees on January 3. b. On January 1 of the current year, Laughter purchases an insurance policy that covers two years, $8,000.
Answer #1
answer
Journal Entry - adjusting entries on 31st december working notes (WN)
particulars debit amount credit amount 1 salary expense for 5 days = 8000
salary a/c dr                           (WN-1) 4800 salary expense per day = 8000/5 1600
      to salary payable a/c 4800
therefore, salary expense for 3 working days
insurance a/c dr                   (WN-2) 4000 =1600*3 4800
      to prepaid insurance a/c 4000
supplies expense a/c dr   (WN-3) 8400 2 insurance policy of 8000 that covers 2 years
      to office supplies a/c 8400 therefore, insurance expense per year = 8000/2
4000
unearned service revenue a/c dr (WN-4) 2600
      to service revenue a/c 2600 3 supplies expense
opening balance 4300
accounts recievables a/c dr 3000 add: purchase 5600
      to service revenue a/c 3000 less: closing balance -1500
8400
depreciation A/c dr (equipment) 3000
      to accumulated depreciation a/c 3000
4 revenue earned for landscape design (40%)
depreciation A/c dr (truck) 2200 =unearned*40%
      to accumulated depreciation a/c 2200 =6500*40%
2600
interest a/c dr 250
      to interest payable a/c 250